38) (20 points) Suppose the demand and supply curves RD 10 2P for a product are given by A. Find the equilibrium price and quantity B. If the current price of the product is $4, what is the quantity demanded and the quantity supplied? How would you describe this situation, and what would you expect to happen in this market? C. If the current price of the product is $1.5, what is the quantity demanded and the quantity supplied? How would you describe this situation, and what would you expect to happen in this market? D. Suppose the demand changes to QD 15-2P. Find the new equilibrium price and quantity and indicate whether the change in demand indicate an increase or a decrease in Demand and why?
Answer #1

A. For equilibrium: Q​​​​​​D = Q​​​​​​S

10 - 2P = 2 + 2P

4P = 8

P = 8 / 4 = $2 (equilibrium price)

Q = 10 - 2(2) = 6 (equilibrium quantity)

B. At price of $4:

Qd = 10 - 2(4) = 2

Qs = 2 + 2(4) = 10

So, there will be an excess supply of 8 units in the market. This will have a downward pressure on price.

C. At price = $1.5

Qd = 10 - 2(1.5) = 7

Qs = 2 + 2(1.5) = 5

So, there will be an excess demand of 2 units in the market. This will have an upward pressure on the price.

D. For new equilibrium

15 - 2P = 2 + 2P

4P = 13

P = 13 / 4 = $3.25 (new equilibrium price)

Q = 15 - 2(3.25) = 8.5 (new equilibrium quantity)

So, the change in the demand indicate an increase in the demand. It is because autonomous consumption increases from 10 to 15.